# HyperAIRI: a plug-and-play algorithm for precise hyperspectral image reconstruction in radio interferometry

### v1

@@ 2.2 Hyperspectral Fourier sampling in RI
where $\nu_{l}$ is the observation frequency, and $c$ is the speed of light. Consequently, for a fixed set of baselines, the $uv$-coverage becomes more concentrated towards the centre of the Fourier plane, i.e. focusing on the low spatial Frequency content, at the lower end of the observation frequency bandwidth. At the higher end of the observation frequency bandwidth, the same sampling pattern dilates radially away from the centre of the Fourier plane, i.e. acquiring the high spatial Frequency content. Figure 1 shows this spectral-frequency-dependent expansion of the $uv$-sampling pattern across three representative channels in a simulated VLA $uv$-coverage. By design, observations at lower-frequency channels are characterised by lower spatial resolution but higher sensitivity, and conversely, higher-frequency channels offer higher resolution but lower sensitivity due to the scaling of the $uv$-coverage.
@@
Figure 1: A set of simulated VLA hyperspectral sampling patterns of different channels. The three panels show the normalised uvuv sampling patterns for the first, middle and last channels of an observation with frequencies 1.00, 1.34 and 1.70 GHz, respectively. The coordinates are in the unit of wavelength.
2.3 RI hyperspectral inverse problem
Focusing on intensity imaging and assuming a narrow field of view without atmospheric or instrumental perturbations, at a given spectral channel $l$, the process of measuring visibilities $\bm{y}_{l}\in\mathbb{C}^{M}$ from the true sky intensity image ${\overline{\bm{x}}}_{l}\in\mathbb{R}^{N}_{+}$ can be modelled by the following discrete linear system:
@@ 2.3 RI hyperspectral inverse problem
where $\bm{\Phi}_{l}\in\mathbb{C}^{M\times N}$ represents the measurement operator at channel $l$, and $\bm{n}_{l}\in\mathbb{C}^{M}$ denotes measurement noise, modelled as a complex white Gaussian noise with mean zero and a standard deviation of $\tau_{l}$. The measurement operator $\bm{\Phi}_{l}$ models incomplete non-uniform Fourier sampling where ${\mathbf{G}_{l}}\in\mathbb{C}^{M\times N^{\prime}}$ is a sparse non-uniform Fourier interpolation matrix, with each row corresponding to a compact convolutional interpolation kernel centred at the corresponding Fourier mode, ${\mathbf{F}}\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$ denotes the 2D Fourier transform, and ${\mathbf{P}}\in\mathbb{R}^{N^{\prime}\times N}$ is a zero-padding operator, also encompassing the correction for the convolutional interpolation kernels (Fessler & Sutton, 2003).
In practice, the noise $\bm{n}_{l}$ may be non-white, in which case a diagonal noise-whitening matrix $\mathbf{\Theta}_{l}\in\mathbb{R}^{M}_{+}$ is incorporated into $\mathbf{\Phi}_{l}$ and applied to the measurements $\bm{y}_{l}$, such that $\bm{\Phi}_{l}=\mathbf{\Theta}_{l}\mathbf{G}_{l}\mathbf{FP}$. Each diagonal element of $\mathbf{\Theta}_{l}$ is the inverse of the noise standard deviation of the corresponding visibility. With this adjustment known as natural weighting, the linear measurement model in (3) remains valid for the whitened visibilities $\bm{y}_{l}$, whereby $\tau_{l}=1$. On a further note, the RI measurement process can be affected by the so-called $w$-effect, a phase modulation arising from the non-coplanarity of the antenna array which is non-negligible in wide-field imaging. In this context, the operator $\bm{\Phi}_{l}$ can account for this effect via additional row-wise convolutions (Dabbech et al., 2017).
@@
3 Optimisation-based hyperspectral imaging
Building on HyperSARA, we introduce its unconstrained variant, Hyper-uSARA, tailored for hyperspectral imaging with unknown noise statistics and calibration errors. This variant is particularly well suited to large-scale real observational data, where flexibility and efficiency are critical. We begin by revisiting the optimisation theory underpinning our method, focusing on the FB for solving unconstrained minimisation problems. We then describe the hyperspectral SARA prior, which promotes joint-channel average sparsity, and present the full algorithmic structure of Hyper-uSARA.
3.1 Optimisation theory
The linear measurement model taking the form (4) is widely used across various imaging tasks, such as denoising, super-resolution, inpainting, image deblurring, and magnetic resonance imaging. While the measurement operator $\Phi(\cdot)$ varies across image modalities, the underlying inverse problem is often ill-posed due to incomplete and noisy measurements. Optimisation methods are thus appealing for their versatility in regularising inverse problems by injecting sophisticated image models.
Within the convex optimisation framework, the inverse problem is addressed by solving a minimisation problem with objective function
$\underset{\mathbf{X}\in\mathbb{R}^{N\times L}}{\min}\;f(\mathbf{X})+\lambda r(\mathbf{X})$ (5)
@@ 3.1 Optimisation theory
When at least one of the two terms is differentiable–in this case the data fidelity term ff–problems of the form (5) can be solved with FB. The iteration structure consists of a gradient descent step (forward) for the differentiable data-fidelity term, followed by and denoising step (backward) via the proximity operator of the non-differentiable regularisation term. At any iteration $i$, the image cube update reads:
@@
The gradient of $f(\cdot)$ reads $\nabla f\left(\mathbf{X}\right)=\mathrm{Re}\{\Phi^{\dagger}\left(\Phi\left(\mathbf{X}\right)\right)\}-\mathrm{Re}\{\Phi^{\dagger}(\mathbf{Y})\}$, where $\Phi^{\dagger}(\cdot)$ is the adjoint of hyperspectral measurement operator and $\mathbf{X}^{\mathrm{dirty}}=\mathrm{Re}\{\Phi^{\dagger}(\mathbf{Y})\}$ is the back-projected data, commonly referred to as the dirty image cube. The parameter $\gamma$ is the step size, upper-bounded by $\|\Phi^{\dagger}(\Phi(\cdot))\|_{\mathrm{S}}22$2In practice, $\gamma$ is typically fixed to $1.98/\|\Phi^{\dagger}(\Phi(\cdot))\|_{\mathrm{S}}$. to ensure the algorithm’s convergence, with $\|\cdot\|_{\mathrm{S}}$ denoting the spectral norm (Bauschke et al., 2017).
By definition, the proximity operator of a regularisation function $r$ is given by
@@
When $r$ is separable across channels, i.e. taking the form $r(\mathbf{X})=\sum_{l=1}^{L}r_{l}(\bm{x}_{l})$, naturally the problem (5) decomposes into $L$ independent monochromatic imaging problems. To exploit spectral correlations and improve reconstruction quality, joint regularisation terms are preferred for hyperspectral RI imaging.
3.2 Hyperspectral SARA prior and Hyper-uSARA
@@
The choice of the prior encoded via the regularisation function $r$ in (5) greatly affects imaging precision. In monochromatic RI imaging, the sparsity-based SARA prior promotes average sparsity of the sought image across a collection of 8 orthonormal wavelet bases and the Dirac basis, denoted as $\bm{\Psi}\in\mathbb{R}^{N\times bN}$ with $b=9$. Its extension to hyperspectral imaging, the HyperSARA prior, is underpinned by the assumption that the image cube is the combination of a few radio sources, each having a distinct spectral behaviour (Abdulaziz et al., 2019). As such, the HyperSARA prior promotes the low-rankness of the image cube, whose rank is upper-bounded by the number of radio sources, and its joint-sparsity over the spectral channels in the SARA dictionary. The former is enforced via a nuclear norm which penalises the $\ell_{1}$-norm of the vector of the singular values of $\mathbf{X}$. The latter is enforced via a $\ell_{2,1}$-norm, defined as the $\ell_{1}$-norm of the vector whose elements are the $\ell_{2}$-norms of the rows of $\bm{\Psi}^{\dagger}\mathbf{X}$. Focusing on intensity imaging, the prior also includes non-negativity of the target image cube. The HyperSARA algorithm takes a constraint minimisation task, involving bounds on the hyperspectral noise cube, and is solved using the PD-FB.
@@
Underpinned by a constrained data fidelity term, HyperSARA requires accurate knowledge of the noise statistics, which can be challenging in practice, with calibration errors often dominating the theoretical noise. To address this, we formally propose the unconstrained formulation of the HyperSARA minimisation task. Furthermore, we discard the low-rankness term, though accepting a slight loss in precision, since it involves singular value decomposition of the image cube estimate at every iteration, a computationally expensive operation for large-scale imaging even when adopting 3D faceting as in Faceted HyperSARA (Thouvenin et al., 2023a, b). This simplification also eliminates the need to balance the average joint-sparsity and low-rankness terms, while reducing the number of regularisation parameters. The resulting unconstrained variant of HyperSARA underpinned by an average joint-sparsity prior is dubbed Hyper-uSARA. The regularisation function reads
$r(\mathbf{X})=\sum^{bN}_{n=1}{\epsilon\log\left(\epsilon^{-1}\|((\bm{\Psi}^{\dagger}\mathbf{X})^{\top})_{n}\|_{2}+1\right)}+\iota_{\mathbb{R}_{+}^{N\times L}}(\mathbf{X}),$ (8)
@@
where $\epsilon$ is a hyper parameter, $(\cdot)_{n}$ represent the $n$-th column of the argument matrix, $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm, and $\iota_{\mathbb{R}_{+}^{N\times L}}(\cdot)$ is an indicator function [^3] The indicator function $\iota_{\mathbb{R}_{+}^{N\times L}}(\mathbf{X})=0$ if all elements in $\mathbf{X}$ are positive, otherwise $\iota_{\mathbb{R}_{+}^{N\times L}}(\mathbf{X})=\infty$. imposing positivity constraint on each pixel. To address the non-convex regularisation function $r$, the minimisation problem of the form (5) is addressed via a re-weighting scheme that iteratively solves a sequence of convex minimisation tasks. More precisely, at each re-weighting iteration a minimisation task involving a surrogate convex regularisation function $\widetilde{r}$ instead of $r$ (Repetti & Wiaux, 2021), such that
@@
The weighted average joint-sparsity regulariser is given by
@@
where $\mathbf{W}\in\mathbb{R}^{bN\times bN}$ is a diagonal matrix, whose diagonal elements $w_{n}$ form the vector $\bm{w}\in\mathbb{R}^{bN}$. The weights $\bm{w}$ are updated at each re-weighting cycle from the solution of the preceding minimisation task as
@@
At each re-weighting iteration $j$, the objective function is updated using the current $\bm{w}_{j}$, and is solved via the FB (6). The Hyper-uSARA algorithm is summarised in Algorithm 1. The proximity operator of the composite regularisation function (9) is computed via a sub-iterative dual-FB (Combettes & Pesquet, 2011) illustrated in Algorithm 2. At a given $\mathbf{Z}\in\mathbb{R}^{N\times L}$, the respective proximity operators for the positivity constraint and the weighted $\ell_{2,1}$-norm sparsity term have closed-form solutions such that
@@
$\left(\left[\mathrm{max}\left\{\|(\mathbf{Z}^{\top})_{n}\|_{2}-\eta w_{n},0\right\}\frac{(\mathbf{Z}^{\top})_{n}}{\|(\mathbf{Z}^{\top})_{n}\|_{2}}\right]_{1\leq n\leq N}\right)^{\top}.$
As in Faceted-HyperSARA, the average joint-sparsity regularisation is applied in parallel to each 3D image facet using the same splitting strategy. Formally, the FB gradient step only requires the dirty image and the application of $\Phi^{\dagger}(\Phi(\cdot))$, keeping memory and computational cost tied to the image size (Vijay Kartik et al., 2017). Hence, Hyper-uSARA can achieve better efficiency for large-scale imaging tasks under tight computational constraints.
Algorithm 1 Hyper-uSARA
1:given $\mathbf{X}^{0}$, $\mathbf{X}^{\mathrm{dirty}}$, $\Phi(\cdot)$, $\Phi^{\dagger}(\cdot)$ // Initialisation
2:given $\gamma$, $\lambda$, $\epsilon$ // Step size, regularisation parameter, noise floor level
3:given $\xi_{1}$, $\xi_{2}$ // Tolerance for re-weighting and FB
4:given $J$, II // Maximum iterations for re-weighting and FB
5:// Re-weighting loop
6:repeat for $j=1,2,\ldots,J$
7:  $\bm{w}^{j}=\epsilon\left/\left(\epsilon+\sqrt{\sum_{l=1}^{L}\left(\bm{\Psi}^{\dagger}\bm{x}^{j-1}_{l}\right)^{2}}\right)\right$.
8:  // FB loop
9:  $\mathbf{X}^{j,0}=\mathbf{X}^{j-1}$
10:  repeat for $i=1,2,\ldots,I$
11:        $\hat{\mathbf{X}}^{j,i}=\mathbf{X}^{j,i-1}-\gamma\mathrm{Re}\left\{\Phi^{\dagger}\left(\Phi\left(\mathbf{X}^{j,i-1}\right)\right)-\mathbf{X}^{\mathrm{dirty}}\right\}$
12:        $\mathbf{X}^{j,i}=\mathrm{prox}_{\gamma\lambda\widetilde{r}}\left(\hat{\mathbf{X}}^{j,i}\right)$ // Using Algorithm 2
[…]
14:  $\mathbf{X}^{j}=\mathbf{X}^{j,i}$
15:until $\left.\left\|\mathbf{X}^{j}-\mathbf{X}^{j-1}\right\|_{\mathrm{F}}\middle/\left\|\mathbf{X}^{j-1}\right\|_{\mathrm{F}}\right.<\xi_{1}$
16:return $\mathbf{X}^{j}$
Algorithm 2 Solving proximity operator of (9) with dual-FB
1:given $\mathbf{X}^{0}$, $\mathbf{\hat{X}}$, $\bm{\Psi}$, $\bm{\Psi}^{\dagger}$, $\bm{w}$ // Initialisation
2:given $\gamma$, $\lambda$ // Step size and regularisation parameter
@@
Concerning the choice of the regularisation parameters involved, both $\lambda$ and $\epsilon$ are inferred from the image domain noise level. Terris et al. (2023) proposed estimating the image domain noise level at each frequency channel as
Under this consideration, and following Thouvenin et al. (2023a), the noise floor level $\epsilon$ in the $\ell_{2}$-norm of the wavelet coefficients is then approximated such that
and the regularisation parameter is chosen as $\lambda=\epsilon/\gamma$ (Thouvenin et al., 2023a).
4 HyperAIRI
In this section, we present the HyperAIRI algorithm, which can be seen as the PnP counterpart of Hyper-uSARA within the PnP-FB. We subsequently detail the architecture of HyperAIRI denoisers and the associated training strategy, and we formalise the algorithmic structure of our proposed approach.

### v2

@@ II.2 Hyperspectral Fourier sampling in RI
where $\nu_{l}$ is the observation frequency, and $c$ is the speed of light. Consequently, for a fixed set of baselines, the $uv$-coverage becomes more concentrated towards the center of the Fourier plane, at the lower end of the observation frequency bandwidth. The same sampling pattern dilates radially away from the center of the Fourier plane, at the higher end of the observation frequency bandwidth. Figure 1 shows this spectral-frequency-dependent dilation of the $uv$-sampling pattern across three representative channels using a simulated VLA $uv$-coverage. By design, observations at lower-frequency channels provides lower spatial resolution but higher sensitivity, and conversely, higher-frequency channels offer higher resolution but lower sensitivity due to the scaling of the $uv$-coverage.
@@
Figure 1: A set of simulated VLA hyperspectral sampling patterns of different channels. The three panels show the normalized uvuv sampling patterns for the first, middle and last channels of an observation with frequencies 1.00, 1.34 and 1.70 GHz, respectively. The coordinates are in the unit of wavelength.
II.3 RI hyperspectral inverse problem
Focusing on intensity imaging and assuming a narrow field of view without atmospheric or instrumental perturbations, at a given spectral channel $l$, the process of measuring visibilities $\bm{y}_{l}\in\mathbb{C}^{M}$ from the true sky intensity image ${\overline{\bm{x}}}_{l}\in\mathbb{R}^{N}_{+}$ can be modeled by the following discrete linear system:
@@ II.3 RI hyperspectral inverse problem
where $\bm{\Phi}_{l}\in\mathbb{C}^{M\times N}$ represents the measurement operator at channel $l$, and $\bm{n}_{l}\in\mathbb{C}^{M}$ denotes measurement noise, modeled as a complex white Gaussian noise with mean zero and a standard deviation of $\tau_{l}$. The measurement operator $\bm{\Phi}_{l}$ models incomplete non-uniform Fourier sampling where ${\mathbf{G}_{l}}\in\mathbb{C}^{M\times N^{\prime}}$ is a sparse non-uniform Fourier interpolation matrix, with each row corresponding to a compact convolutional interpolation kernel centered at the corresponding Fourier mode, denotes the 2D Fourier transform, and ${\mathbf{P}}\in\mathbb{R}^{N^{\prime}\times N}$ is a zero-padding operator, also encompassing the correction for the convolutional interpolation kernels (J. A. Fessler & B. P. Sutton, 2003).
In practice, the noise $\bm{n}_{l}$ may be non-white, in which case a diagonal noise-whitening matrix $\mathbf{\Theta}_{l}\in\mathbb{R}^{M}_{+}$ is incorporated into $\mathbf{\Phi}_{l}$ and applied to the measurements $\bm{y}_{l}$, such that $\bm{\Phi}_{l}=\mathbf{\Theta}_{l}\mathbf{G}_{l}\mathbf{FP}$. Each diagonal element of $\mathbf{\Theta}_{l}$ is the inverse of the noise standard deviation of the corresponding visibility. With this adjustment known as natural weighting, the linear measurement model in (3) remains valid for the whitened visibilities $\bm{y}_{l}$, whereby $\tau_{l}=1$. On a further note, the RI measurement process can be affected by the so-called $w$-effect, a phase modulation arising from the non-coplanarity of the antenna array which is non-negligible in wide-field imaging. In this context, the operator $\bm{\Phi}_{l}$ can account for this effect via additional row-wise convolutions (A. Dabbech et al., 2017).
@@
III Optimization-based hyperspectral imaging
Building on HyperSARA, we introduce its unconstrained variant, Hyper-uSARA, tailored for hyperspectral imaging with unknown noise statistics and calibration errors. This variant is particularly well suited to large-scale real observational data, where flexibility and efficiency are critical. We begin by revisiting the optimization theory underpinning our method, focusing on the FB for solving unconstrained minimization problems. We then describe the hyperspectral SARA prior, which promotes joint-channel average sparsity, and present the full algorithmic structure of Hyper-uSARA.
III.1 Optimization theory
The linear measurement model taking the form (4) is widely used across various imaging tasks, such as denoising, super-resolution, inpainting, image deblurring, and magnetic resonance imaging. While the measurement operator $\Phi(\cdot)$ varies across image modalities, the underlying inverse problem is often ill-posed due to incomplete and noisy measurements. Optimization methods are thus appealing for their versatility in regularizing inverse problems by injecting sophisticated image models.
Within the convex optimization framework, the inverse problem is addressed by solving a minimization problem with objective function
$\underset{\mathbf{X}\in\mathbb{R}^{N\times L}}{\min}\;f(\mathbf{X})+\lambda r(\mathbf{X}),$ (5)
@@ III.1 Optimization theory
When at least one of the two terms is differentiable–in this case the data fidelity term ff–problems of the form (5) can be solved with FB. The iteration structure consists of a gradient descent step (forward) for the differentiable data-fidelity term, followed by and denoising step (backward) via the proximity operator of the non-differentiable regularization term. At any iteration $i$, the image cube update reads
@@
The gradient of $f(\cdot)$ reads $\nabla f\left(\mathbf{X}\right)=\mathrm{Re}\{\Phi^{\dagger}\left(\Phi\left(\mathbf{X}\right)\right)\}-\mathrm{Re}\{\Phi^{\dagger}(\mathbf{Y})\}$, where $\Phi^{\dagger}(\cdot)$ is the adjoint of hyperspectral measurement operator and $\mathbf{X}^{\mathrm{dirty}}=\mathrm{Re}\{\Phi^{\dagger}(\mathbf{Y})\}$ is the back-projected data, commonly referred to as the dirty image cube. The parameter $\gamma$ is the step size, upper-bounded by $2/\|\Phi^{\dagger}(\Phi(\cdot))\|_{\mathrm{S}}$ to ensure the algorithm’s convergence [^2] In practice, $\gamma$ is typically fixed to $1.98/\|\Phi^{\dagger}(\Phi(\cdot))\|_{\mathrm{S}}$., with $\|\cdot\|_{\mathrm{S}}$ denoting the spectral norm (H. H. Bauschke et al., 2017).
By definition, the proximity operator of a regularization function $r$ is given by
@@
When $r$ is separable across channels, i.e. taking the form $r(\mathbf{X})=\sum_{l=1}^{L}r_{l}(\bm{x}_{l})$, naturally the problem (5) decomposes into $L$ independent monochromatic imaging problems. To exploit spectral correlations and improve reconstruction quality, joint regularization terms are preferred for hyperspectral RI imaging.
III.2 Hyperspectral SARA prior and Hyper-uSARA
@@
The choice of the prior encoded via the regularization function $r$ in (5) greatly affects imaging precision. In monochromatic RI imaging, the sparsity-based SARA prior promotes average sparsity of the sought image across a collection of 8 orthonormal wavelet bases and the Dirac basis, denoted as $\bm{\Psi}\in\mathbb{R}^{N\times bN}$ with $b=9$. Its extension to hyperspectral imaging, the HyperSARA prior, is underpinned by the assumption that the image cube is the combination of a few radio sources, each having a distinct spectral behavior (A. Abdulaziz et al., 2019). As such, the HyperSARA prior promotes the low-rankness of the image cube, whose rank is upper-bounded by the number of radio sources, and its joint-sparsity over the spectral channels in the SARA dictionary. The former is enforced via a nuclear norm which penalizes the $\ell_{1}$-norm of the vector of the singular values of $\mathbf{X}$. The latter is enforced via an $\ell_{2,1}$-norm, defined as the $\ell_{1}$-norm of the vector whose elements are the $\ell_{2}$-norms of the rows of $\bm{\Psi}^{\dagger}\mathbf{X}$. Focusing on intensity imaging, the prior also includes non-negativity of the target image cube. The HyperSARA algorithm takes a constraint minimization task, involving bounds on the hyperspectral noise cube, and is solved using the PD-FB.
@@
Underpinned by a constrained data fidelity term, HyperSARA requires accurate knowledge of the noise statistics, which can be challenging in practice, with calibration errors often dominating the theoretical noise. To address this, we formally propose the unconstrained formulation of the HyperSARA minimization task. Furthermore, we discard the low-rankness term, though accepting a slight loss in precision, since it involves singular value decomposition of the image cube estimate at every iteration, a computationally expensive operation for large-scale imaging even when adopting 3D faceting as in Faceted HyperSARA (P.-A. Thouvenin et al., 2023a, b). This simplification also eliminates the need to balance the average joint-sparsity and low-rankness terms, while reducing the number of regularization parameters. The resulting unconstrained variant of HyperSARA underpinned by an average joint-sparsity prior is dubbed Hyper-uSARA. The regularization function reads
$r(\mathbf{X})=$ $\sum^{bN}_{n=1}{\epsilon\log\left(\epsilon^{-1}\|((\bm{\Psi}^{\dagger}\mathbf{X})^{\top})_{n}\|_{2}+1\right)}$ (8)
$\qquad+\iota_{\mathbb{R}_{+}^{N\times L}}(\mathbf{X}),$
@@
where $\epsilon$ is a hyperparameter, $(\cdot)_{n}$ represent the $n$-th column of the argument matrix, $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm, and $\iota_{\mathbb{R}_{+}^{N\times L}}(\cdot)$ is an indicator function [^3] The indicator function $\iota_{\mathbb{R}_{+}^{N\times L}}(\mathbf{X})=0$ if all elements in $\mathbf{X}$ are positive, otherwise $\iota_{\mathbb{R}_{+}^{N\times L}}(\mathbf{X})=\infty$. imposing positivity constraint on each pixel. To address the non-convex regularization function $r$, the minimization problem of the form (5) is addressed via a re-weighting scheme that iteratively solves a sequence of convex minimization tasks. More precisely, at each re-weighting iteration a minimization task involving a surrogate convex regularization function $\widetilde{r}$ instead of $r$ (A. Repetti & Y. Wiaux, 2021), such that
@@
The weighted average joint-sparsity regularizer is given by
@@
where $\mathbf{W}\in\mathbb{R}^{bN\times bN}$ is a diagonal matrix, whose diagonal elements $w_{n}$ form the vector $\bm{w}\in\mathbb{R}^{bN}$. The weights $\bm{w}$ are updated at each re-weighting cycle from the solution of the preceding minimization task as
@@
At each re-weighting iteration $j$, the objective function is updated using the current $\bm{w}_{j}$, and is solved via the FB (6). The Hyper-uSARA algorithm is summarized in Algorithm III.2. The proximity operator of the composite regularization function (9) is computed via a sub-iterative dual-FB (P. L. Combettes & J.-C. Pesquet, 2011) illustrated in Algorithm III.2. At a given $\mathbf{Z}\in\mathbb{R}^{N\times L}$, the respective proximity operators for the positivity constraint and the weighted $\ell_{2,1}$-norm sparsity term have closed-form solutions such that
@@
$\mathrm{prox}_{\eta\|\cdot\|_{1,2,\bm{w}}}(\mathbf{Z})$ $=\Biggl(\Biggl[\mathrm{max}\!\left\{\|(\mathbf{Z}^{\top})_{n}\|_{2}-\eta w_{n},0\right\}$ (13)
$\qquad\qquad\frac{(\mathbf{Z}^{\top})_{n}}{\|(\mathbf{Z}^{\top})_{n}\|_{2}}\Biggr]_{1\leq n\leq N}\Biggr)^{\!\top}.$
As in Faceted-HyperSARA, the average joint-sparsity regularization is applied in parallel to each 3D image facet using the same splitting strategy. Formally, the FB gradient step only requires the dirty image and the application of $\Phi^{\dagger}(\Phi(\cdot))$, keeping memory and computational cost tied to the image size (S. Vijay Kartik et al., 2017). Hence, Hyper-uSARA can achieve better efficiency for large-scale imaging tasks under tight computational constraints.
Concerning the choice of the regularization parameters involved, both $\lambda$ and $\epsilon$ are inferred from the image-domain noise level. M. Terris et al. (2023) proposed estimating the image-domain noise level at each frequency channel as
Under this consideration, and following P.-A. Thouvenin et al. (2023a), the noise floor level $\epsilon$ in the $\ell_{2}$-norm of the wavelet coefficients is then approximated such that
and the regularization parameter is chosen as $\lambda=\epsilon/\gamma$ (P.-A. Thouvenin et al., 2023a).
Algorithm 1. Hyper-uSARA
1:// Initialization
2:given $\mathbf{X}^{0}$, $\mathbf{X}^{\mathrm{dirty}}$, $\Phi(\cdot)$, $\Phi^{\dagger}(\cdot)$
3:// Step size, regularization parameter, noise floor level
4:given $\gamma$, $\lambda$, $\epsilon$
5:// Tolerance for re-weighting and FB
6:given $\xi_{1}$, $\xi_{2}$
7:// Maximum iterations for re-weighting and FB
8:given $J$, II
9:// Re-weighting loop
10:repeat for $j=1,2,\ldots,J$
11:  $\bm{w}^{j}=\epsilon\left/\left(\epsilon+\sqrt{\sum_{l=1}^{L}\left(\bm{\Psi}^{\dagger}\bm{x}^{j-1}_{l}\right)^{2}}\right)\right$.
12:  // FB loop
13:  $\mathbf{X}^{j,0}=\mathbf{X}^{j-1}$
14:  repeat for $i=1,2,\ldots,I$
15:        $\hat{\mathbf{X}}^{j,i}=\mathbf{X}^{j,i-1}-\gamma\mathrm{Re}\left\{\Phi^{\dagger}\left(\Phi\left(\mathbf{X}^{j,i-1}\right)\right)-\mathbf{X}^{\mathrm{dirty}}\right\}$
16:        $\mathbf{X}^{j,i}=\mathrm{prox}_{\gamma\lambda\widetilde{r}}\left(\hat{\mathbf{X}}^{j,i}\right)$ // Using Algorithm III.2
@@
18:  $\mathbf{X}^{j}=\mathbf{X}^{j,i}$
19:until $\left.\left\|\mathbf{X}^{j}-\mathbf{X}^{j-1}\right\|_{\mathrm{F}}\middle/\left\|\mathbf{X}^{j-1}\right\|_{\mathrm{F}}\right.<\xi_{1}$
20:return $\mathbf{X}^{j}$
Algorithm 2. Solving proximity operator of (9) with dual-FB
1:given $\mathbf{X}^{0}$, $\mathbf{\hat{X}}$, $\bm{\Psi}$, $\bm{\Psi}^{\dagger}$, $\bm{w}$ // Initialization
2:given $\gamma$, $\lambda$ // Step size and regularization parameter
@@
IV HyperAIRI
In this section, we present the HyperAIRI algorithm, which can be seen as the PnP counterpart of Hyper-uSARA within the PnP-FB. We subsequently detail the architecture of HyperAIRI denoisers and the associated training strategy, and we formalize the algorithmic structure of our proposed approach.